OpenStudy (anonymous):

Can someone explain the rule of Negative Fractional exponents, for example: 1 - y^-5

OpenStudy (anonymous):

negative exponents. to change to positive ones, if it is up bring it down, if it is down bring it up. \[\frac{1}{y^{-5}}=\frac{y^5}{1}=y^5\]

OpenStudy (amistre64):

fractional exponents?

OpenStudy (anonymous):

meant negative exponents with fractions i think from the example

OpenStudy (amistre64):

your example is good in that case; but it could have been a sympathy medal :)

OpenStudy (anonymous):

yeah probably. like how i wrote \[\frac{y^5}{1}\]?